OpenStudy (anonymous):

how do i find the 3rd root of 40 ?

OpenStudy (anonymous):

do you mean the cube root?

OpenStudy (anonymous):

it's like , an exponent of 3 outside of the square root of 40

OpenStudy (anonymous):

That is the cube root\[\sqrt[3]{?}\]

OpenStudy (anonymous):

yea like that

OpenStudy (anonymous):

\[\sqrt[3]{40}=\sqrt[3]{2*2*2*5}=2\sqrt[3]{5}\]

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

it can be called the 3rd root also swirling just like the 2nd root or 4th root or 5th... i believe it is better known as the cube root

jhonyy9 (jhonyy9):

\[\sqrt[3]{40} = \sqrt[3]{2^{3}*5}=2\sqrt[3]{5}\]